# Bottom-up Policy Optimization: Your Language Model Policy Secretly Contains Internal Policies

### v1

@@ 3.2 Internal Policy Entropy Dynamics
How do Internal modules influence the residual stream? To further understand how internal modules shape the residual stream in Qwen models with progressive reasoning pattern, we analyze residual cosine similarity, which quantifies how each module writes to the residual pathway (hu2025affects). For a given layer $l$, we compute $\text{cossim}(\mathbf{A}^{l},\mathbf{H}^{l-1})$ for self-attention and $\text{cossim}(\mathbf{F}^{l},\mathbf{H}^{l-1}+\mathbf{A}^{l})$ for the FFN. A cosine similarity near zero indicates writing new, orthogonal features; negative values indicate feature suppression; and positive values indicate amplification of existing features.
The FFN modulates the residual stream in a stage-dependent manner across all models: in lower layers, it injects largely orthogonal features to support exploration; in middle layers, it suppresses vague signals while integrating parametric knowledge in FFN, corresponding to the Integration stage; and in upper layers, it amplifies andintegrates features to drive convergence. Across all models, the final layer exhibits a sharp directional shift, underscoring its critical role in final prediction (gupta2025llms; agarwal2025unreasonable).
Takeaway Build on the definition of internal policy and corresponding entropy, we reveal the following findings: Llama models exhibit limited exploration in both self-attention and FFN modules, leading to constrained exploration in internal reasoning. Qwen-series display a staged, progressive reasoning pattern, with the FFN following an Exploration–Integration–Convergence process. Qwen3 shows a more stable FFN reasoning structure than Qwen2.5, and stronger attention amplification to the residual stream, enabling richer information gain.
4 Internal Policy Optimization
Building on our previous analysis, we observe that reasoning emerges progressively, with different internal policy regions serving distinct roles and exhibiting varied behaviors. This raises a key question: Can we approach optimization from a bottom-up perspective to align with this progressive emergence? To address this, we extend GRPO which originally designed for language model policy optimization to directly optimize these internal policies:
$\frac{1}{G}\sum^{G}_{i=1}\frac{1}{|\mathbf{o}_{i}|}\sum_{t=1}^{|\mathbf{o_{i}}|}\left\{\min\left[\hat{r}_{i,t}\hat{A}_{i,t},\text{clip}(\hat{r}_{i,t},1-\epsilon,1+\epsilon)\hat{A}_{i,t}\right]\right\}$
where we still sample from $\pi_{\theta_{\text{old}}}$, $\hat{r}_{i,t}=\frac{\pi_{\text{Layer}}^{l}(o_{i,t}|\mathbf{q},\mathbf{o}_{i,<t})}{\pi^{l}_{\text{Layer,old}}(o_{i,t}|\mathbf{q},\mathbf{o}_{i,<t})}$ is the modified importance ratio for current optimizing policy $\pi^{l}_{\text{Layer}}$ and keep other hyperparameters same with E.q. 4. We present the implementation details in Appendix A.5.
Training setup. Based on the findings in Sec. 3, we select Qwen3-4B (yang2025qwen3) (non-thinking mode) for investigation. For the training set, we randomly sample 5k entries from DeepMath-103k (he2025deepmath). We train the models using the verl framework (sheng2025hybridflow). The prompt batch size is 128, with 8 rollouts generated per prompt. The sampling temperature during training is set to 1.0, and the maximum context length is set to 9,216 tokens. We update the model with a mini-batch size of 32 for 300 steps and a learning rate of 1e-6. For Qwen3-4B, the region boundaries identified in Sec. 3.2 lie at layers 6 and 26. Additionally, we focus on aligning the internal policy of the penultimate layer, which is critical for the final prediction. Accordingly, we compare the internal policies $\pi_{\text{Layer}}^{6}$, $\pi_{\text{Layer}}^{26}$, and $\pi_{\text{Layer}}^{35}$ optimized via InterGRPO, against the overall policy $\pi_{\theta}$ optimized with GRPO (he2025deepmath).
@@ 4 Internal Policy Optimization
Different training dynamics of internal policy. As shown in Figure 11, both internal layer policies collapse after 50 training steps, suggesting that optimizing internal policies alone is fundamentally misaligned with the objective of the language model policy $\pi_{\theta}$. Nevertheless, distinct patterns emerge. The entropy of $\pi_{\theta}$ rapidly converges to zero under standard RL training (cui2025entropy), a well-known phenomenon encountered by various methods (peng2025simko; zhu2025surprising; dong2025agentic). For the penultimate layer policy $\pi_{\text{Layer}}^{35}$, entropy shows minor fluctuations before aligning with $\pi_{\theta}$. But it suffers from severe repetition causing excessively long responses, likely because final decision-making is confined to the last layer (gupta2025llms). In contrast, the knowledge integration layer policy $\pi_{\text{Layer}}^{26}$ exhibits unstable and increasing entropy, also resulting in longer outputs. The lower-layer internal policy $\pi_{\text{Layer}}^{6}$ maintains more stable entropy growth, with response lengths converging closer to those of $\pi_{\theta}$.
@@
Analysis of internal policy optimization. Given the distinct training dynamics, we further investigate the underlying mechanism of internal policy optimization. Surprisingly, we find that this process induces a significant feature refinement in the model’s internal states. Taking layer 6 $\pi_{\text{Layer}}^{6}$ for illustration, Figure 12 (a) shows that as optimization progresses, the similarity between its hidden states $\mathbf{H}^{6}$ and the final layer representations increases. This suggests that internal optimization forces the bottom layer to preemptively capture high-level reasoning information, providing a more robust foundation for subsequent internal reasoning. While the entropy change dynamics in Figure 12(b) indicate that the optimized internal policy progressively converges, the PPL trajectory in Figure 12(c) reveals a clear trade-off: excessive alignment leads to performance collapse, suggesting that only a limited number of optimization steps are optimal.
Table 1: Avg@$K$ results on MATH500, AMC23, AIME24 and AIME25. Bold and unserlined numbers denote the best an second-best results for each $K$. We select the last FFN layer that exhibits a positive exploration signal (i.e., $\Delta H_{\text{FFN}}^{l}>0$) as $\pi_{\text{Layer}}^{l}$.
@@
     BuPO ($\pi_{\text{Layer}}^{6}\rightarrow\pi_{\theta},s_{\text{inter}}=30$) 81.09+4.21 84.90+2.49 36.88+4.69 31.15+2.30 58.51+3.43
@@
     BuPO ($\pi_{\text{Layer}}^{6}\rightarrow\pi_{\theta},s_{\text{inter}}=20$) 89.22+3.28 87.76 54.06+4.58 34.38+0.76 66.36+2.13
@@
     BuPO ($\pi_{\text{Layer}}^{27}\rightarrow\pi_{\theta},s_{\text{inter}}=20$) 27.50+0.00 49.79+3.72 0.63+0.00 0.42+0.32 19.59+1.01
@@
     BuPO ($\pi_{\text{Layer}}^{31}\rightarrow\pi_{\theta},s_{\text{inter}}=20$) 37.66+2.82 62.05+5.16 4.69+2.19 6.77+4.58 27.79+3.68
@@
5 Bottom-up Policy Optimization
All prior RL methods for LLMs optimize the language model policy in a holistic manner (ouyang2022training; rafailov2023direct; guo2025deepseek). In Sec. 3, we decompose the final output $\mathbf{H}^{L}$ into an intermediate representation $\mathbf{H}^{l}$ and the subsequent residual contribution $\mathbf{S}^{l+1}$. This decomposition suggests that aligning the internal policy $\pi_{\text{Layer}}^{l}$ associated with $\mathbf{H}^{l}$ may facilitate alignment of the overall policy $\pi_{\theta}$. We empirically support this intuition in Sec. 4, where we optimize the internal policy alone and observe pronounced feature refinement in lower layers. These findings further motivate a bottom-up alignment strategy: by aligning finer-grained internal policies first, we explore whether the overall policy can be guided to reason more effectively.
To this end, we propose Bottom-up Policy Optimization (BuPO), which sequentially optimizes the internal layer policies $\pi_{\text{Layer}}^{l}$ followed by the language model policy $\pi_{\theta}$. The overall training objective is:
where $s_{\text{cur}}$ denotes the current training step, and $s_{\text{inter}}$ specifies the number of training steps of the internal layer policy. We present the detailed training algorithm in Appendix A.1.
Training setup. Specifically, we focus on the Qwen3 series, which maintain a stable Exploration-Integration-Convergence pattern, using Qwen3-4B and Qwen3-8B. For comparison, we select Llama-OctoThinker-3B-Base and Llama-OctoThinker-8B-Base from the Llama series, as these models demonstrate improved RL training performance after additional mid-training based on Llama-3.2-Base (wang2025octothinker). The maximum context length for Llama is set to 4096 tokens. We select the last FFN layer that exhibits a positive exploration signal (i.e., $\Delta H^{l}_{\text{FFN}}>0$) as $\pi_{\text{Layer}}^{l}$. Other RL training settings follow Sec. 4 and summarized in Table 5.
[…]
@@ 5.1 Main Results
We report the Avg@k results in Table 1. Surprisingly, adopting a bottom-up optimization perspective in BuPO leads to consistent improvements over RL baseline algorithms across benchmarks and models, achieving superior average performance. On the Qwen3-4B model, BuPO yields gains of 4.58 points on AIME24 and 0.76 points on AIME25 compared to GRPO. Similarly, the Qwen-8B model shows improvements of 3.65 points on AIME24 and 2.40 points on AIME25. The Llama-series models exhibit a similar optimization trend under BuPO, achieving an average improvement of 1.01 points on Llama-OctoThinker-3B-Base and 3.68 points on Llama-OctoThinker-8B-Base. Overall, these results suggest that aligning finer-grained internal policies in the early stages of training can effectively guide the language model policy toward improved reasoning.
For a more comprehensive evaluation, we report the averaged Pass@K results of BuPO with GRPO on MATH500, AMC23, AIME24, and AIME25, with $K$ ranging from 1 to 256. As shown in Figure 13, BuPO consistently achieves a favorable trade-off across a wide range of $K$ values. On Qwen3-8B, BuPO attains the best performance for all $K$ values, while on Qwen3-4B, the only exception occurs at $K=256$. For all Llama-series models, BuPO achieves the best results across all $K$ values, yielding gains of 7.48 points on Llama-OctoThinker-3B-Base and 7.93 points on Llama-OctoThinker-8B-Base in averaged Pass@256. These results further indicate that bottom-up optimization effectively enhances the reasoning capacity of the language model policy.
@@ 5.2 Analysis
We further visualize the training dynamics of BuPO following the settings in Sec. 5.1. As shown in Figure 17(a), by training the internal layer policy at an early stage, all models exhibit enhanced entropy exploration initially. For Qwen models, optimizing the first region boundary layer 6 maintains stable exploration, consistent with Figure 11. For the Llama-series, we observe a consistent increase during the bottom alignment stage, indicating that feature refinement in lower layers effectively provides a larger exploration space for language model policy.
@@
In Sec. 4, we investigate the optimization of the internal layer policy. The remarkable effectiveness of this approach further validates our bottom-up policy optimization strategy. First, early stage internal optimization compels the bottom layers to preemptively capture high-level reasoning information, establishing a better foundation for subsequent learning (Figure 12 (a,b)). Second, excessive training steps lead to model collapse, as the mismatch between the rollout policy and the optimized policy results in increased PPL (Figure 12 (c)). We conclude that moderate bottom optimization of the internal policy can further boost the learning capability of the base model, making BuPO a practical RL method.
@@
5.3 Ablation Study
Ablation of Bottom Optimization Steps.
Sec. 4 reveals that as the optimization steps of the internal layer policy progress, the language model policy eventually collapses. We further conduct an ablation study of how $s_{\text{inter}}$ affects the performance of BuPO. As shown in the upper panel of Table 2, as the bottom alignment steps increase, the performance of the language model policy drops dramatically. This aligns with the finding in Figure 12 and further supports our conclusion in Sec. 5.2 that only moderate bottom optimization can boost overall policy learning effectively.
Ablation of optimized internal policy.
We further investigated the impact of the optimized internal policy following Sec.4, with a focus on the region boundaries highlighted in our takeaways. As shown in the lower panel of Table 2, with fixed bottom optimization steps, the bottom-up optimization strategy achieves superior performance compared to GRPO. The entropy dynamics in Figure 17(a) indicate that the alignment of $\pi_{\text{Layer}}^{26}$ induces an entropy spike, leading to significant exploration expansion, aligning with its strong performance. The selection and investigation of internal policies remain promising directions for future research.
6 Related Work
6.1 Reinforcement Learning with Verifiable Rewards
Reinforcement learning (RL) has proven effective for enhancing large language models (LLMs), most notably through Reinforcement Learning from Human Feedback (RLHF), which aligns model outputs with human preferences (sutton1998reinforcement; rafailov2023direct; jaech2024openai; hu2025reinforce++; wang2025adaptive). More recently, Reinforcement Learning with Verifiable Rewards (RLVR) has gained traction for its ability to foster reasoning in LLMs using rule-based rewards (li2025attention; peng2025simko; tan2025zero). Despite these advances, prior work has primarily focused on optimizing the overall language model policy, overlooking the rich information latent within its internal residual streams. In this work, we shift the focus inward to analyze these streams, constructing internal policies that reveal consistent and distinct reasoning structures across different models. Building on this finding, we introduce a bottom-up optimization strategy that directly optimizes internal policies at intermediate layers, strengthening the underlying reasoning structure and leading to improved overall performance.
6.2 Interpretability of Black-box LLMs
Interpretability tools mitigate the opacity of black-box LLMs by revealing their internal logic. A substantial body of prior work investigates how LLMs reason and memory (yu2023neuron; tan-etal-2025-neural; lindsey2025biology; gupta2025llms; hu2025affects), with particular emphasis on the self-attention module (zhou2024role; jin2025massive) and the feed-forward network (FFN) module (dai2022knowledge; meng2022locating; geva2023dissecting). Insights into these internal mechanisms provide a new perspective for algorithmic optimization (li2025attention; liu2025attention). In this work, we conduct a systematic analysis of internal hidden states from a policy-centric perspective and identify various internal reasoning patterns across model families. We further observe that reasoning emerges progressively from lower to higher layers. Motivated by this finding, we propose bottom-up policy optimization, which leads to improved reinforcement learning algorithms.
@@ 7 Conclusion
In this work, we propose a novel decomposition of language model policies into Internal Layer Policies and Internal Modular Policies, enabling fine-grained analysis of how reasoning evolves across layers and modules. Through systematic entropy analysis, we uncover distinct reasoning patterns: early layers maintain high entropy for exploration, while top layers converge to near-zero entropy for refinement. Notably, we identify architectural differences—Llama models exhibit abrupt convergence in final layers, whereas Qwen-series models, especially Qwen3, demonstrate progressive, human-like reasoning with FFNs playing a crucial structural role. Our exploration of internal policy optimization reveals a key insight: optimizing internal policies encourages lower layers to encode high-level reasoning signals early, reconstructing foundational reasoning capabilities rather than merely adjusting surface-level outputs. Motivated by this, we introduce Bottom-up Policy Optimization (BuPO), a reinforcement learning paradigm that directly targets internal layer policy during early training. Extensive experiments on complex reasoning benchmarks demonstrate that BuPO consistently improves learning efficiency and achieves superior Avg@KK/Pass@$K$ performance, validating the effectiveness of layer-aware policy optimization.
Limitations.
We believe this work can be further extended. While our experiments focus on the commonly used Qwen and Llama model families, extending the study to additional architectures would help assess the broader applicability of our approach. Moreover, our analysis is conducted under specific experimental settings, and different RLVR configurations may exhibit varying behaviors. Finally, the selection of bottom-up optimized internal layers and training steps may depend on model characteristics and could be adjusted to achieve optimal performance in different settings.
Future directions.
Future directions involve adopting advanced interpretability tools to facilitate both the understanding of internal mechanisms and algorithmic design. Further exploration of the internal reasoning patterns in Llama, Qwen, and other model families will feedback into model development. Moreover, a comprehensive understanding of internal policies can make the bottom-up perspective more generalizable.

### v2

@@ 4 Internal Policy Optimization
Different Training Dynamics of Internal Policy.
@@
As shown in Figure 8, both internal layer policies collapse after 50 training steps. Nevertheless, distinct patterns emerge. For the penultimate layer policy $\pi_{\text{Layer}}^{35}$, entropy shows minor fluctuations before aligning with $\pi_{\theta}$. But it suffers from repetition causing excessively long responses, likely because final decision-making is confined to the last layer (Gupta et al., 2025). In contrast, the last integration region policy $\pi_{\text{Layer}}^{26}$ exhibits unstable and increased entropy. And the last exploration region policy $\pi_{\text{Layer}}^{6}$ maintains stable entropy growth, with response lengths converging closer to $\pi_{\theta}$.
@@
Analysis of Internal Policy Optimization. We further investigate the underlying mechanism of internal policy optimization. Surprisingly, we find that this process induces a significant feature refinement in the model’s internal states. Taking $\pi_{\text{Layer}}^{6}$ for illustration, Figure 9 (a) shows that as optimization progresses, the similarity between its hidden states $\mathbf{H}^{6}$ and the final layer representations increases. This suggests that internal optimization forces the bottom layer to preemptively capture high-level reasoning information, providing a more robust foundation for subsequent internal reasoning. While the entropy change dynamics in Figure 9(b) indicate that the optimized internal policy progressively converges, the PPL trend in Figure 9(c) reveals a clear trade-off: excessive alignment leads to performance collapse, suggesting limited steps of optimization are optimal.
Table 1: Avg@$K$ results on MATH500, AMC23, AIME24 and AIME25. Bold and underlined denote the best and second best.
@@
     BuPO 81.09+4.21 84.90+2.49 36.88+4.69 31.15+2.30 58.51+3.43
@@
     BuPO 89.22+3.28 87.76 54.06+4.58 34.38+0.76 66.36+2.13
@@
     BuPO 27.50+0.00 49.79+3.72 0.63+0.00 0.42+0.32 19.59+1.01
@@
     BuPO 37.66+2.82 62.05+5.16 4.69+2.19 6.77+4.58 27.79+3.68
5 Bottom-up Policy Optimization
All prior RL methods for LLMs optimize the language model policy in a holistic manner (Ouyang et al., 2022; Rafailov et al., 2023; Guo et al., 2025). In Sec. 3, we decompose the final output $\mathbf{H}^{L}$ into an intermediate representation $\mathbf{H}^{l}$ and the subsequent residual contribution $\mathbf{S}^{l+1}$. This decomposition suggests that aligning the internal policy $\pi_{\text{Layer}}^{l}$ associated with $\mathbf{H}^{l}$ may facilitate alignment of the overall policy $\pi_{\theta}$. We empirically support this intuition in Sec. 4, where we optimize the internal policy alone and observe pronounced feature refinement in lower layers. These findings further motivate a bottom-up alignment strategy: by aligning fine-grained internal policies first, we explore whether the overall policy can be guided to reason more effectively.
To this end, we propose Bottom-up Policy Optimization (BuPO), which sequentially optimizes the internal layer policies $\pi_{\text{Layer}}^{l}$ followed by the language model policy $\pi_{\theta}$:
where $s_{\text{cur}}$ denotes the current training step, and $s_{\text{inter}}$ specifies the number of training steps of the internal layer policy.
@@ 5 Bottom-up Policy Optimization
Training Setup. Using Qwen (Yang et al., 2025) and Llama (Grattafiori et al., 2024; Wang et al., 2025b) backbones, we select the last FFN layer with a positive exploration signal ($\Delta H^{l}_{\text{FFN}}>0$) as $\pi_{\text{Layer}}^{l}$. This assigns $\pi_{\text{Layer}}^{6}$ for both Qwen3-4B and Qwen3-8B, $\pi_{\text{Layer}}^{27}$ for Llama-OctoThinker-3B-Base, and $\pi_{\text{Layer}}^{31}$ for Llama-OctoThinker-8B-Base. The training steps for internal layer policy are set to $s_{\text{inter}}=30$ for Qwen3-4B and $s_{\text{inter}}=20$ for the other LLMs. Refer to Appendix A.1 and A.6 for detailed algorithm and setup.
Evaluation Setup. We evaluate BuPO against several RL baselines, including GRPO, PPO (Sutton et al., 1998), Reinforce++ (Hu, 2025) and RLOO (Ahmadian et al., 2024). The benchmarks cover : MATH (Lightman et al., 2023), AMC23 (MAA, 2023), AIME24, and AIME25 (MAA, 2024, 2025). Due to high output variance in reasoning tasks, we report Avg@$K$ (Pass@1 averaged over $K$ outputs). For AIME24/25, we set $K=32$, and for others $K=16$. Additionally, we evaluate an unbiased Pass@$K$ metric with $K$ up to 256 to obtain comprehensive evaluation. Detailed evaluation setups are shown in Appendix A.6.
@@ 5.1 Main Results
As shown in Table  1, adopting the bottom-up perspective in BuPO leads to consistent Avg@$K$ improvements over RL baseline algorithms across benchmarks and models, achieving superior average performance. On Qwen3-4B, BuPO yields gains of 4.69 points on AIME24 and 2.30 points on AIME25 compared to GRPO. Similarly, the Qwen-8B model shows improvements of 4.58 points on AIME24 and 0.76 points on AIME25. The Llama series models exhibit a similar optimization trend under BuPO, achieving an average improvement of 1.01 points on Llama-OctoThinker-3B-Base and 3.68 points on Llama-OctoThinker-8B-Base. Overall, these results suggest that aligning fine-grained internal policies in the early stages of training can effectively guide the language model policy toward improved reasoning.
For a comprehensive evaluation, we report the averaged Pass@$K$ results of BuPO with GRPO, with $K$ ranging from 1 to 256. As shown in Figure 10, BuPO consistently achieves a favorable trade-off across a wide range of $K$. On Qwen3-8B, BuPO attains the best performance for all $K$ values, while on Qwen3-4B, the only exception occurs at $K=256$. For all Llama models, BuPO achieves the best results across all $K$ values, yielding gains of 7.48 points on Llama-OctoThinker-3B-Base and 7.93 points on Llama-OctoThinker-8B-Base in averaged Pass@256. These results further indicate the internal policy optimization effectively enhances the reasoning capacity of $\pi_{\theta}$. We provide detailed Pass@$K$ performance in Figure 18.
@@ 5.2 Analysis
We further visualize the training dynamics of BuPO. As shown in Figure 14(a), by training the internal layer policy at an early stage, all models exhibit enhanced entropy exploration initially. For Qwen models, optimizing the first region boundary layer 6 maintains stable exploration. For Llama, we observe increased entropy during the bottom alignment stage, indicating that feature refinement in lower layers effectively provides a larger exploration space for language model policy.
@@
In Sec. 4, we investigate the direct optimization of the internal layer policy. The remarkable effectiveness of this approach further validates our designed policy optimization strategy from bottom-up perspective. First, internal policy optimization at early stage compels the bottom layers to preemptively capture high-level reasoning information, establishing a better foundation for subsequent learning (Figure 9 (a,b)). Second, excessive training steps lead to model collapse, as the mismatch between the rollout policy and the optimized policy results in increased PPL (Figure 9 (c)). We conclude that moderate bottom optimization of the internal policy can further boost the learning capability of the overall language model policy, making BuPO a practical RL method.
Ablation Study.
(1) Effect of Bottom Optimization Steps. We further conduct an ablation study of how $s_{\text{inter}}$ affects the performance of BuPO. As shown in the upper panel of Table 2, as the bottom alignment steps increase, the performance of the language model policy drops dramatically. This aligns with the finding in Figure 9 and further supports our conclusion in Sec. 5.2 that only moderate bottom optimization can boost overall policy learning effectively. (2) Effect of Optimized Internal Policy. We further investigated the impact of the optimized internal policy, focus on the region boundaries in Sec. 3.2. As shown in the lower panel of Table 2, BuPO variants employing different $\pi_{\text{Layer}}^{l}$ exhibit superior performance over GRPO with a fixed $s_{\text{inter}}$. The entropy dynamics in Figure 14(a) indicate that the alignment of $\pi_{\text{Layer}}^{26}$ induces an entropy spike, leading to significant exploration expansion, aligning with its strong performance. The selection and investigation of internal policies remain promising directions for future research.
6 Related Work
Reinforcement Learning with Verifiable Rewards.
Reinforcement learning (RL) has proven effective for enhancing large language models (LLMs)  (Sutton et al., 1998; Rafailov et al., 2023; Jaech et al., 2024; Hu, 2025; Wang et al., 2025a). More recently, Reinforcement Learning with Verifiable Rewards (RLVR) has gained traction for its ability to foster reasoning in LLMs using rule-based rewards (Li et al., 2025; Peng et al., 2025; Tan et al., 2025b). In this work, we shift the focus to internal policies, revealing consistent yet distinct reasoning structures across models. Building on these observations, we propose bottom-up policy optimization, which directly optimizes internal layer policies in at early training stages. This targeted optimization refines internal reasoning representations and ultimately leads to improved overall performance.
@@ 6 Related Work
Interpretability of Black-box LLMs.
Interpretability tools mitigate the opacity of black-box LLMs by revealing their internal logic. A substantial body of prior work investigates how LLMs reason and memory (Yu and Ananiadou, 2024; Tan et al., 2025a; Lindsey et al., 2025; Gupta et al., 2025; Hu et al., 2025b), with particular emphasis on the self-attention (Zhou et al., 2025; Jin et al., 2025) and the feed-forward network (Dai et al., 2022; Meng et al., 2022; Geva et al., 2023). Insights into these internal mechanisms provide a new perspective for algorithmic optimization (Li et al., 2025; Liu et al., 2025). In this work, we conduct a systematic analysis of internal hidden states from a policy-centric perspective and identify various internal reasoning patterns across models. Motivated by these observations, we propose bottom-up policy optimization, which refines internal features and leads to a effective RL algorithm.
@@ 7 Conclusion
In this paper, we first propose a novel decomposition of language model policies into internal layer policies and internal modular policies, revealing systematic reasoning patterns via entropy analysis. We observe a universal transition from high-entropy exploration in early layers to deterministic convergence in higher layers, with Qwen models exhibiting progressive, human-like reasoning, in contrast to the abrupt convergence of Llama models. Motivated by these insights, we propose Bottom-up Policy Optimization (BuPO), a reinforcement learning paradigm that explicitly optimizes internal layer policies at early training stages. Extensive experiments on complex reasoning benchmarks demonstrate the effectiveness of BuPO. Through in-depth analysis, we show that aligning internal layer policies reconstructs foundational reasoning capabilities, thereby further enhancing the reasoning capacity of entire language model policy.
Impact Statement
This work aims to advance the fields of Reinforcement Learning (RL) and Mechanistic Interpretability by investigating the internal reasoning patterns of Large Language Models (LLMs). Our research contributes to the transparency of black-box LLMs by decomposing the language model policy into Internal Layer and Modular policies. By revealing how internal reasoning emerges progressively (as observed in Qwen LLMs) or abruptly (as in Llama LLMs), our work helps understand model internal mechanisms and enhances algorithmic efficiency, a crucial step toward building more controllable AI systems. Furthermore, our proposed algorithm, BuPO, improves the performance of LLMs on complex reasoning tasks, potentially enhancing the utility of AI in educational and scientific domains.